HETAV PATEL

Grade 11,

A And B Are Two Vectors and R Is The Resultant. For What Angle Is IT Maximum For What Angle Is It Minimum

A And B Are Two Vectors and R Is The Resultant.
  1. For What Angle Is IT Maximum
  2. For What Angle Is It Minimum

Grade:11

4 Answers

Manas Shukla
7 years ago
R is maximum when angle between them is 0 degrees since they just get added.
R is minimum when vectors cancel out the effects of one another
angle between them being 180 degrees
Piyush Kumar Behera
7 years ago
@Manas Shukla how can you say that at \theta=180 they cancel out each other ,It is not given that both are equal ,so the term “cancel out each other” doesnot suit.However it can be say that the net resaultant is decreased due to opposing vectors.
Piyush Kumar Behera
7 years ago
Here is another solution:
 
We know the expression for R=(A2+B2+2ABcos\theta)½
As A& B2 are positive and constant so the value of R will be Maximum when cos\theta is Maximum.
Hence cos\theta is maximum at \theta=0o so R is maximum at \theta=0o.
Similarly cos\theta is minimum at \theta=180o so R is minimum at \theta=180o.
If you are helped by my solution then please approve the solution to encourage me.
pradyumna choudhury
7 years ago
for angle to be maximum cos\varnothing should be maximum 
so as we know cos\varnothing is maximum at 0 degree 
cos\o is minimum at 180 degree